Consider the initial value problem y + 2y – 15y = 0, y0) = a, y0 = 1 Find the value of a so that the solution to the init
Answer #1

& Griven the Irp y+2y-15y=0, y(o)=d, g!) = 1 (0²+2D-15)y=0 where is a Auxiliary qualion m? +2m -15=0 ) (mts) (m-3) = 0 => m